OpenStudy (anonymous):

How to factor this equation w^2+18w+77

OpenStudy (anonymous):

w^2+11w+7w+77 from here solve it

OpenStudy (anonymous):

you need two numbers which when multiplied together give 77 and when add give 18 these are 7 and 11 so factors are (x + 7)(x + 11)

OpenStudy (anonymous):

w^2+11w+7x+77 = w(w+11)+7(w+11) (w+7)(w+11)

OpenStudy (anonymous):

either x+7 = 0 or x+11 = 0 so roots are -7,-11

OpenStudy (anonymous):

oh sorry - replace x by w!!!

OpenStudy (anonymous):

soooo its neg instead of pos . ?

OpenStudy (anonymous):

the roots of the equation (w+7)(w+11) = 0 are negative as w + 7 = 0 subtract 7 from both sides w +7 - 7 = 0-7 w = -7 same thing for w+11

OpenStudy (anonymous):

you called (w+7)(w+11) an equation - its not - its a function of w (w+7)(w+11) = 0 is an equation

OpenStudy (anonymous):

in order to solve it you factor it first